# Cocompact Fuchsian groups with a modular embedding

### v2

@@ Abstract
A Fuchsian group $\Gamma$ has a modular embedding if its adjoint trace field is a totally real number field and every unbounded Galois conjugate $\Gamma^{\sigma}$ comes equipped with a holomorphic (or conjugate holomorphic) map ${\phi^{\sigma}:\mathbb{B}^{1}\to\mathbb{B}^{1}}$ intertwining the actions of $\Gamma$ and $\Gamma^{\sigma}$ on the Poincaré disk $\mathbb{B}^{1}$. This paper provides the first cocompact nonarithmetic Fuchsian groups with a modular embedding that are not commensurable with a triangle group. The main result, proved using period domains, is that any immersed totally geodesic complex curve on a complex hyperbolic $2$-orbifold has a modular embedding. Another consequence is that there are infinitely many signatures for which no finite-volume quotient of $\mathbb{B}^{1}$ with that signature can be an immersed totally geodesic curve on a complex hyperbolic $2$-orbifold.
@@ 1 Introduction
This gives a negative answer to [20, Prob. 1] in the cocompact case. The examples are immersed totally geodesic curves on nonarithmetic complex hyperbolic $2$-orbifolds, i.e., finite-volume quotients of the unit ball $\mathbb{B}^{2}$ in $\mathbb{C}^{2}$ equipped with its Bergman metric of constant holomorphic sectional curvature $-1$ and holomorphic isometry group $\operatorname{PU}(2,1)$. Their modular embedding comes from the following very general result.
@@
It is worth noting that the nonarithmetic orbifolds constructed by Deligne and Mostow [8, 18] may well be inadequate for proving 1.1. While Cohen and Wolfart proved that Deligne–Mostow lattices also have the analogue of a modular embedding [7], all known immersed totally geodesic complex curves on nonarithmetic Deligne–Mostow orbifolds are commensurable with triangle orbifolds. In fact, since nonarithmetic complex hyperbolic $2$-orbifolds have only finitely many immersed totally geodesic curves [2, 3], it may be the case that all totally geodesic curves on Deligne–Mostow orbifolds are related to triangle orbifolds. The examples in this paper arise from the lattices constructed by Deraux, Parker, and Paupert [10].
1.2 also implies arithmeticity of reflection groups associated with certain properly immersed totally geodesic suborbifolds of finite-volume complex hyperbolic $2$-orbifolds. Let $\mathcal{Q}[n_{1},\dots,n_{4}]$ be a quadrilateral in $\mathbb{B}^{1}$ with (cyclically ordered) interior angles $\pi/n_{j}$ and $\widehat{\Delta}[n_{1},\dots,n_{4}]<\operatorname{Isom}(\mathbb{B}^{1})$ be the group generated by reflections in the sides of $\mathcal{Q}[n_{1},\dots,n_{4}]$. Ricker [19], following up on work of Schmutz Schaller and Wolfart [20], proved for $[n_{1},\dots,n_{4}]$ any of $[2,2,2,t]$, $[2,2,t,t]$, or $[2,t,2,t]$ with $t\in\{3,4,6\}$ that the orientation-preserving index two subgroup $\Delta[n_{1},\dots,n_{4}]$ of $\widehat{\Delta}[n_{1},\dots,n_{4}]$ has a modular embedding if and only if $\widehat{\Delta}[n_{1},\dots,n_{4}]$ is an arithmetic reflection group. Having a modular embedding is a commensurability invariant, so 1.2 and [19] combine to imply the following.
@@
Suppose $[n_{1},\dots,n_{4}]$ is $[2,2,2,t]$, $[2,2,t,t]$, or $[2,t,2,t]$ with $t\in\{3,4,6\}$ and $Z$ is a quotient of $\mathbb{B}^{1}$ commensurable with $\widehat{\Delta}[n_{1},\dots,n_{4}]\backslash\mathbb{B}^{1}$, where $\widehat{\Delta}[n_{1},\dots,n_{4}]$ is the group generated by reflections in the sides of a quadrilateral of angles $\pi/n_{j}$. If $Z$ can be holomorphically immersed as a totally geodesic curve on any complex hyperbolic $2$-manifold, then $\widehat{\Delta}[n_{1},\dots,n_{4}]$ is an arithmetic reflection group.
@@ Acknowledgements
Thanks are due to Curtis McMullen for his interest in the paper and for detailed comments on an early draft and to Greg Baldi for several conversations related to this paper and the relevant literature. This material is based upon work supported by Grant Number DMS-2203555 from the National Science Foundation.
@@ 2 Preliminaries
A finite-volume complex hyperbolic $2$-orbifold is then a quotient $\Lambda\backslash X$ by a lattice $\Lambda<G$. Properly holomorphically immersed totally geodesic subspaces
$\Gamma\backslash\mathbb{B}^{1}\looparrowright\Lambda\backslash X$
@@
Totally geodesic curves also have a dynamical interpretation. Indeed, they correspond precisely with the closed orbits for the action of $W$ on the homogeneous space $\Lambda\backslash G$. Specifically, if $[g]$ denotes the point $\Lambda g\in\Lambda\backslash G$, then $[g]W$ is a closed orbit if and only if $\Gamma_{g}$ is a lattice in $gWg^{-1}$. See [2, §8.1] for more on dynamical characterizations of geodesic subspaces.
@@
For all problems considered in this paper, there is no loss of generality in passing to a subgroup of finite index, so $\Lambda$ is henceforth assumed to be torsion-free. The adjoint action of $\Lambda$ defines a $k$-algebraic group $\mathbf{G}$ containing $\Lambda$ as a Zariski dense subgroup, and let $\widehat{\mathbf{G}}$ be the restriction of scalars of $\mathbf{G}$ from $k$ to $\mathbb{Q}$. After possibly conjugating and passing to a further subgroup of finite index, $\Lambda$ is then a subgroup of $\widehat{\mathbf{G}}(\mathbb{Z})$ where the integral structure is defined through the polarized integral variation of Hodge structure $\widehat{\mathbb{V}}$ on $\Lambda\backslash X$ that arises from lifting $\Lambda$ to $\operatorname{SU}(2,1)$ and then taking the natural action on $\mathbb{C}^{3}$; see [3, Thm. 1.3.1].
Every simple factor of $\widehat{\mathbf{G}}(\mathbb{R})$ is of Hodge type (e.g., see [2, Thm. 1.5]), which implies in this case that
@@
that is equivariant for the inclusion of $\Lambda$ in $\widehat{\mathbf{G}}(\mathbb{Z})$. If
@@
As manifolds, the possible target domains $D_{\sigma}$ are in one-to-one correspondence with parabolic subgroups $Q<G_{\sigma}(\mathbb{C})\cong\operatorname{PSL}_{3}(\mathbb{C})$ so that $Q\cap G_{\sigma}$ is compact and $G_{\sigma}/(Q\cap G_{\sigma})$ is an open $G_{\sigma}$-orbit in $\operatorname{PSL}_{3}(\mathbb{C})/Q$. If $Q$ is the stabilizer of a line (or, dually, a plane), then $Q\cap G_{\sigma}$ is conjugate to $K$ and $D_{\sigma}$ is simply $X$ with its unique complex structure (up to complex conjugation). If $Q$ is the minimal parabolic subgroup of $\operatorname{SL}_{3}(\mathbb{C})$, namely the stabilizer of a full flag, then $V_{\sigma}\cong L$ is the maximal compact subgroup of $W$. This discussion is recorded in the following lemma, where the reader should consult [12, p. 169-170] or [5, §16.4] for details.
@@
Retain all notation from Section 2, supposing moreover that $\Lambda<G$ is a torsion-free nonarithmetic lattice with $[g]W\subset\Lambda\backslash G$ the closed $W$-orbit associated with a totally geodesic curve on the complex hyperbolic $2$-manifold $\Lambda\backslash X$. Let $\Gamma$ be the associated finite coarea Fuchsian group, so there is a proper holomorphic totally geodesic immersion
$\Gamma\backslash\mathbb{B}^{1}\looparrowright\Lambda\backslash X$
@@
It is noted in [9] that the Fuchsian group $\Gamma$ from 4.1 is nonarithmetic, so it suffices by 1.2 to prove that $\Gamma$ is not commensurable with a triangle group. Let $d$ denote the potential index of $\Gamma$ in the triangle group. A simple computer check using orbifold Euler characteristics and the following facts:
@@
Figure 1: Quadrilaterals with vertices at the fixed points of $a$, $b$, $c$, and $abc$

### v3

@@ Abstract
A Fuchsian group $\Gamma$ has a modular embedding if its adjoint trace field is a totally real number field and every unbounded Galois conjugate $\Gamma^{\sigma}$ comes equipped with a holomorphic (or conjugate holomorphic) map $\phi^{\sigma}:\mathbb{B}^{1}\to\mathbb{B}^{1}$} intertwining the actions of $\Gamma$ and $\Gamma^{\sigma}$ on the Poincaré disk $\mathbb{B}^{1}$. This paper provides the first cocompact nonarithmetic Fuchsian groups with a modular embedding that are not commensurable with a triangle group. The main result, proved using period domains, is that any immersed totally geodesic complex curve on a complex hyperbolic $2$-orbifold has a modular embedding. Another consequence is arithmeticity of totally geodesic curves on finite-volume complex hyperbolic surfaces that are commensurable with quotients of $\mathbb{B}^{1}$ by the group generated by reflections in quadrilaterals satisfying certain angle conditions.
@@ 1 Introduction
This gives a negative answer to [20, Prob. 1] in the cocompact case. The examples are immersed totally geodesic curves on nonarithmetic complex hyperbolic $2$-orbifolds, i.e., finite-volume quotients of the unit ball $\mathbb{B}^{2}$ in $\mathbb{C}^{2}$ equipped with its Bergman metric of constant holomorphic sectional curvature 1-1 and holomorphic isometry group $\operatorname{PU}(2,1)$. Their modular embedding comes from the following very general result.
@@
It is worth noting that the nonarithmetic orbifolds that were constructed by Deligne and Mostow [8, 18] may well be inadequate for proving 1.1. While Cohen and Wolfart proved that Deligne–Mostow lattices also have the analogue of a modular embedding [7], all known immersed totally geodesic complex curves on nonarithmetic Deligne–Mostow orbifolds are commensurable with triangle orbifolds. In fact, since nonarithmetic complex hyperbolic $2$-orbifolds have only finitely many immersed totally geodesic curves [2, 3], it may be the case that all totally geodesic curves on Deligne–Mostow orbifolds are related to triangle orbifolds. The examples in this paper arise from the lattices constructed by Deraux, Parker, and Paupert [10].
1.2 also implies arithmeticity of reflection groups associated with certain properly immersed totally geodesic suborbifolds of finite-volume complex hyperbolic $2$-orbifolds. Let $\mathcal{Q}[n_{1},\dots,n_{4}]$ be a quadrilateral in $\mathbb{B}^{1}$ with (cyclically ordered) interior angles $\pi/n_{j}$ and $\widehat{\Delta}[n_{1},\dots,n_{4}]<\operatorname{Isom}(\mathbb{B}^{1})$ be the group generated by reflections in the sides of $\mathcal{Q}[n_{1},\dots,n_{4}]$. Ricker [19], following up on work of Schmutz Schaller and Wolfart [20], proved for $n_{1},\dots,n_{4}]$ any of $[2,2,2,t]$, $[2,2,t,t]$, or $[2,t,2,t]$ with $t\in\{3,4,6\}$ that the orientation-preserving index two subgroup $\Delta[n_{1},\dots,n_{4}]$ of $\widehat{\Delta}[n_{1},\dots,n_{4}]$ has a modular embedding if and only if $\widehat{\Delta}[n_{1},\dots,n_{4}]$ is an arithmetic reflection group. Having a modular embedding is a commensurability invariant, so 1.2 and [19] combine to imply the following.
@@
Suppose $n_{1},\dots,n_{4}]$ is $[2,2,2,t]$, $[2,2,t,t]$, or $[2,t,2,t]$ with $t\in\{3,4,6\}$ and $Z$ is a quotient of $\mathbb{B}^{1}$ commensurable with $\widehat{\Delta}[n_{1},\dots,n_{4}]\backslash\mathbb{B}^{1}$, where $\widehat{\Delta}[n_{1},\dots,n_{4}]$ is the group generated by reflections in the sides of a quadrilateral of angles $\pi/n_{j}$. If $Z$ can be holomorphically immersed as a totally geodesic curve on any complex hyperbolic $2$-manifold, then $\widehat{\Delta}[n_{1},\dots,n_{4}]$ is an arithmetic reflection group.
@@ Acknowledgements
Thanks are due to Curtis McMullen for his interest in the paper and for detailed comments on an early draft, Greg Baldi for several conversations related to this paper and the relevant literature, Joshua Lam for noticing a misstatement in an earlier version, and the referee for suggestions. The author acknowledges support from the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and LabEx CARMIN (ANR-10-LABx-59-01). This material is based upon work supported by Grants DMS-2203555 and DMS-2506896 from the National Science Foundation and the Simons Foundation [SFI-MPS-TSM-00014184, MS].
@@ 2 Preliminaries
A finite-volume complex hyperbolic $2$-orbifold is then a quotient $\X\Lambda\backslash$ X by a lattice $\Lambda<G$. Properly holomorphically immersed totally geodesic subspaces
$\X\Gamma\backslash\mathbb{B}^{1}\looparrowright\Lambda\backslash$ X
@@
Totally geodesic curves also have a dynamical interpretation. Indeed, they correspond precisely with the closed orbits for the action of $W$ on the homogeneous space $\G\Lambda\backslash$ G. Specifically, if $[g]$ denotes the point $\G\Lambda g\in\Lambda\backslash$ G, then $[g]W$ is a closed orbit if and only if $\Gamma_{g}$ is a lattice in $gWg^{-1}$. See [2, §8.1] for more on dynamical characterizations of geodesic subspaces.
@@
[…]
@@
[…]
@@
As manifolds, the possible target domains $D_{\sigma}$ are in one-to-one correspondence with parabolic subgroups $Q<G_{\sigma}(\mathbb{C})\cong\operatorname{PSL}_{3}(\mathbb{C})$ so that $Q\cap G_{\sigma}$ is compact and $G_{\sigma}/(Q\cap G_{\sigma})$ is an open $G_{\sigma}$-orbit in $\operatorname{PSL}_{3}(\mathbb{C})/Q$. If $Q$ is the stabilizer of a line (or, dually, a plane), then $Q\cap G_{\sigma}$ is conjugate to $K$ and $D_{\sigma}$ is simply $X$ with its unique complex structure (up to complex conjugation). If $Q$ is the minimal parabolic subgroup of $\operatorname{SL}_{3}(\mathbb{C})$, namely the stabilizer of a full flag, then $V_{\sigma}\cong L$ is the maximal compact subgroup of $W$. This discussion is recorded in the following lemma, where the reader should consult either [12, p. 169-170] or [5, §16.4] for full details.
@@
Retain all notation from Section 2, supposing moreover that $\Lambda<G$ is a torsion-free nonarithmetic lattice with $\G[g]W\subset\Lambda\backslash$ G the closed $W$-orbit associated with a totally geodesic curve on the complex hyperbolic $2$-manifold $\X\Lambda\backslash$ X. Let $\Gamma$ be the associated finite coarea Fuchsian group, so there is a proper holomorphic totally geodesic immersion
$\X\Gamma\backslash\mathbb{B}^{1}\looparrowright\Lambda\backslash$ X
@@
It is noted in [9] that the Fuchsian group $\Gamma$ from 4.1 is nonarithmetic, so it suffices by 1.2 to prove that $\Gamma$ is not commensurable with a triangle group. Since any lattice in $\operatorname{PU}(1,1)$ containing a triangle group is itself a triangle group [21, §6], it suffices to show that $\Gamma$ is not contained in a triangle group. Let $d$ denote the potential index of $\Gamma$ in the triangle group. A simple computer check using orbifold Euler characteristics and the following facts:
@@
Figure 1: Quadrilaterals with vertices at the fixed points of $a$, $b$, $c$, and abcabc